Use the equation: h = 5t^2 +vt + c
h= HEIGHT
v= VELOCITY
t= TIME
c= distance thrown from

A rock fall off a cliff 80 meters from the ground. How long does it take for the rock to hit the ground?

Answer by Alan3354(69443)

Use the equation: h = 5t^2 +vt + c
h= HEIGHT
v= VELOCITY
t= TIME
c= distance thrown from

A rock fall off a cliff 80 meters from the ground. How long does it take for the rock to hit the ground?
-------------
Heights are measured from the ground up.
If something falls, its height decreases.
--------
Use the equation: h = -5t^2 +vt + c (-5, not +5)
h(t) = -5t^2 + 0t + 80
It hits the ground when h(t) = 0
h(t) = -5t^2 + 0t + 80 = 0
5t^2 = 80
t^2 = 16
t = 4 seconds